OpenStudy (anonymous):

Your startup company Fabulous Fudge Inc. enters a new market. Your marketing consultants estimate the daily demand for your chocolate bars as q(p) = 2000; if p < 1; 1000(-p^2 + 2p + 1); if 1 <= p < 2; 1000(p^2 - 6p + 9); if 2 <= p < 3; 0; if 3 <= p: Here q denotes the quantity demanded per day and p the price per chocolate bar in dollars. (a) Sketch the graph of demand q(p) as a function of price p. (b) Compute the price elasticity of demand E(p). (c) Determine the range of prices p for which demand is elastic, inelastic and unit elastic. (d) If the price per chocolate bar is

OpenStudy (anonymous):

General form of a quadratic: ax² + bx + c if a > 0 (parabola opens up) if a < 0 (parabola opens down)

OpenStudy (anonymous):

Okayyy...

OpenStudy (anonymous):

Can you help me with the equation?

OpenStudy (anonymous):

did you understand the graph ?

OpenStudy (anonymous):

yes i think. you need the demand equation i think

OpenStudy (anonymous):

assume q as y and p as x y is constantly 2000 for any values of x below 1. y is a quadratic equation when x is between 1 and 3. The profiles of the equations are determined by the process I gave in my first reply After that y is constantly 0 for any values of x now which equation do you want

OpenStudy (anonymous):

ok i am calculating

OpenStudy (anonymous):

Your startup company Fabulous Fudge Inc. enters a new market. Your marketing consultants estimate the daily demand for your chocolate bars as q(p) = 2000; if p < 1; 1000(-p^2 + 2p + 1); if 1 <= p < 2; 1000(p^2 - 6p + 9); if 2 <= p < 3; 0; if 3 <= p: Here q denotes the quantity demanded per day and p the price per chocolate bar in dollars. (a) Sketch the graph of demand q(p) as a function of price p. (b) Compute the price elasticity of demand E(p). (c) Determine the range of prices p for which demand is elastic, inelastic and unit elastic. (d) If the price per chocolate bar is 1:2$ and you wish to increase your revenue, should you increase or decrease the unit price? (e) Determine the price(s) p at which the instantaneous rate of change of revenue with respect to price is zero.

OpenStudy (anonymous):

this is the full question. sorry it got cut out

OpenStudy (anonymous):

will it be the second or the third equation?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so.....

OpenStudy (anonymous):

actually I was busy ,,it will be third

OpenStudy (anonymous):

I will return after 10 min

OpenStudy (anonymous):

i will wait here

OpenStudy (anonymous):

are you back???

OpenStudy (anonymous):

hy

OpenStudy (anonymous):

the price for 1 bar is 2$, then you have to take the third equation

OpenStudy (anonymous):

1000(p^2 - 6p + 9); if 2 <= p < 3;

OpenStudy (anonymous):